OpenStudy (anonymous):

need help with this equation 4^2x=32?

OpenStudy (callisto):

~~~~~~~~~~~~~LS~~~~~~~~~~~~~~~~~ \[4^{2x} \]\[= (2^2)^{2x}\]\[=2^{4x}\] ~~~~~~~~~~~~~RS~~~~~~~~~~~~~~~~~ \[32\]\[=2^5\] ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ So, \[4^{2x} = 32\]\[2^{4x}=2^5\]Can you solve it now?

OpenStudy (anonymous):

so it'll be 4x=5 right then that'll be 4/5

OpenStudy (anonymous):

@Callisto